OpenStudy (anonymous):

How do I find the line that is parallel to Y= -3 and goes thru the poine (2,5)

OpenStudy (shadowfiend):

Are you familiar with the formula y = mx + b?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So -3=mx=B?

OpenStudy (shadowfiend):

-3 = mx + b, right.

OpenStudy (shadowfiend):

So is there an x on the left?

OpenStudy (shadowfiend):

Actually, let's try this instead. An equation is usually y = mx + b. This equation is y = -3. Is there an x in that?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Nope its just Y=-3 throught the point (2,5)

OpenStudy (shadowfiend):

Heh. Right. So, there's no slope. What does the line y = -3 look like?

OpenStudy (anonymous):

Straight up and down

OpenStudy (anonymous):

Okay thank you

OpenStudy (shadowfiend):

So you got it? :)

OpenStudy (anonymous):

Yep thanks ** dunno y but there is a serious delay in our responses and then when it pops up its at like 9min

OpenStudy (shadowfiend):

Hm. Strange. We'll look into it, thanks!

OpenStudy (anonymous):

what i am cunfuesd